23 January 2016

04:50 MODS1 collection of transform data
  • prep MODS1 by running transformInit.pro in ~modseng/modsScripts/Support/
  • transform_collect using list (created today):
    • /home/lbto/idl/sext/transform/lists/Stone_D_mods.list
      • 12 stars, PA=0
    • starts with r 10 (11 is BG... we don't really need this for MODS).
      • I started earlier, but using a different list Stone_D_MODS in /home/lbto/idl/wfsc...) and aborted to change the list (hence the images 8,9).
      • clouds came - lost guide star number 18 (image 13 was OK, but transform_collect crashed on image 14...)
    • Stone_D_mods_pa90.list
      • 11 stars, PA=90, star 34 on-axis

05:45 Doug encouraged taking data for MODS2 transform.
  • Edited PCSInstrument.conf file to uncomment the line with the corrected right rotator ZP: 26.95
  • DGH bounced PCS
  • transform_collect
    • Stone_E_mods_1.list (in /home/lbto/idl/sext/transform/lists)
    • Clouds came in and we could not find the guide star. transform_collect crashed.

06:05
  • grpixflats_mods2.cal
    • 15 (one of the 5 UG5 flats) is bad
  • slitflats_mods2.cal (nimgs=1)
    • a move fault. I have seen this several times within the past week on MODS2:
>> slitmask LS5x60x0.6
** ERROR: SLITMASK SLITMASK=3 Move Fault, position at end of move 3 but requested position 5
** Abort, Retry, or Ignore? > r
  • 07:00 --- check exposure times for slitflats because seq issue is triggered by slightly high counts. Need to stay below 30k on blue channel.
  • 07:18 execMODS2 slitflats_mods2.cal (in ~/Service_Observing/SciPrep/MODS2)
    • starts with b 26 and r 15.
      • ND1.5 b 28 is bad (seq issue ND1.5).
      • UG5: b 29 (seq issue first UG5). b30 is OK
DThompson taking over...

### UT 20160123 Engineering
luci?.20160123.NNNN.fits

### LUCI Binocular AGw transform data collection
# Stone F_TransformCollect.xml

Both frames 0001 5x2.75s dark

Manually reconfigure: N30_FieldStop, Br G, N375

Both frames 0002 5x2.75s with N30_FieldStop in FPU, Br G

09:26 Started script Stone F_TransformCollect.xml
ZD 45 @ start, FWHM ~0.45" on both sides
Files 0003 - 0032 on both sides
LUCI1 seems a bit out of focus, LUCI2 too but not as bad
GS11 is a binary, the catalog position may be off
GS45 is fairly faint, probably won't work if seeing is bad

10:21 Script completed! No problems or errors!

Manually execute 45"E sync binocular offset with the telescope service gui...OK

Both frames 0033 offset position for sky subtraction

LUCI seems to be treating sync binocular observation requests to also mean
synchronize the readouts. Noticed when L side took an extra cycle to collimate

### LUCI Binocular pointing and collimation model data collection?

10:46 Pointing logging turned on in PCS
/lbt/data/share/PointingLogs/20160123_SX.log
/lbt/data/share/PointingLogs/20160123_DX.log

Using iraf to send the presets to the WT10/ACT stars
/home/LBTO/POINTING/Pointingstars5 to find the targets

Need to wait for collimation at each position

Left side collimating very slowly. MV was not on. DGH
cleared active optics and we let it re-converge...
Better, but still slow

File# Cell Star ID G* All are 5x2.75s LIR:INT
0034 0 0 0 dark frame
0035 1 ACT0352 4
0036 3 ACT0397 5 cell#2 was too low
0037 6 ACT0184 4 cell#4 no star cell#5 too low
Failed...clouds getting too thick to see guide stars
Blind filter in FW2 inserted

Still heavy clouds, calling the night at 12:57. Not enough
time to make a dent in the pointing/collimation data.

Both LUCIs configured with both blind filters. No mask in
the FPU (for the LUCI2 mask exchange today).


-- %USERSIG{OlgaKuhn - 2016-01-23}%

Comments

 
Topic revision: r3 - 23 Jan 2016, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback